On January 1, 2018, Morey, Inc., exchanged $191,175 for 25 percent of Amsterdam Corporation. Morey appropriately applied the equity method to this investment. At January 1, the book values of Amsterdam’s assets and liabilities approximated their fair values.

On June 30, 2018, Morey paid $640,500 for an additional 70 percent of Amsterdam, thus increasing its overall ownership to 95 percent. The price paid for the 70 percent acquisition was proportionate to Amsterdam’s total fair value. At June 30, the carrying amounts of Amsterdam’s assets and liabilities approximated their fair values. Any remaining excess fair value was attributed to goodwill.

Amsterdam reports the following amounts at December 31, 2018 (credit balances shown in parentheses):

Revenues

$

(298,000

)

Expenses

173,000

Retained earnings, January 1

(272,300

)

Dividends declared, October 1

30,000

Common stock

(500,000

)

Amsterdam’s revenue and expenses were distributed evenly throughout the year and no changes in Amsterdam’s stock have occurred.

a.     Using the acquisition method, calculate the acquisition-date fair value of Amsterdam to be included in Morey's June 30 consolidated financial statements.

b.     Using the acquisition method, calculate the revaluation gain (or loss) reported by Morey for its 25 percent investment in Amsterdam on June 30.

c.      Using the acquisition method, calculate the amount of goodwill recognized by Morey on its December 31 balance sheet (assume no impairments have been recognized).

d.     Using the acquisition method, calculate the noncontrolling interest amount reported by Morey on its June 30 and December 31 consolidated balance sheet.

a.

Acquisition-date fair value

       ???          

b.

???

???

c.

Goodwill

???

d.

Noncontrolling interest on June 30

???

Noncontrolling interest on December 31

???
Answer #1

a) Acquisition-date fair value of Amsterdam to be included in Morey's June 30 consolidated financial statements.

= Amount paid on the date of acquisition/% of stake which is purchased

= 640,500/70%

= 915,000

b) Revaluation gain:-

Investment on 01/01 is $ 191,175 for 25% till 06/30

So, gain for first 6 months is $ 15,625 ((298000-173000)/2*25%). So, book value of investment is $ 206,800 (191,175 + 15,625)

Now, fair value of investment on 06/30 is

= 915,000(a)*25% (stake till 06/30)

= $ 228,750

So, revaluation gain

= Fair value of investment - Book value of investment

= 228,750 - 206,800

= $ 21,950

c) Goodwill on 12/31:-

Fair value of investment as 06/30 = $ 915,000

Book value of investment at 06/30 = Retained earnings as on 01/01 + Profit of 6 months + Common stock

= 272,300 + (298000-173000)/2 + 500,000

= 272,300+62,500+500,000

= 834,800

So, Goodwill = Fair value - Book value

= 915,000-834,800

= $ 80,200

d) (i) Non controlling interest at 06/30

5% of fair value on 06/30. 5% because now 95% control is with Morey, only 5% control is with Amsterdam.

= 915,000*5%

= 45,750

(ii) Non controlling interest at 12/31

= NCI at 06/30 + NCI profit from 06/30 till 12/31- NCI dividend

= 45,750 + ((125000/2)*5%) - (30000*5%)

= 45,750 + 3,125 - 1,500

= 47,375

So, below are the answers in summary:-

a 915,000
b 21,950
c 80,200
d

(i) 45,750

(ii) 47,375